OpenStudy (anonymous):

Which equation can pair with x + 2y = 5 to create an inconsistent system? 2x + 4y = 3 5x + 2y = 3 6x + 12y = 30 3x + 4y = 8

OpenStudy (anonymous):

x+2y=5 3x-15=-6y Solve the 1st for "x", as follows: x=5-2y Substitute that into #2 to get; 3(5-2y)-15=-6y 15-6y-15=-6y

OpenStudy (anonymous):

so is it c ?